Find the exact value:(without Calculator) Please show me how, i have an exam coming up! thank you in advance!
a) 2sin*2pie/3)-4tan*3pie/4
b) 4tan*5pie/3+7cos*(-pie/4)

Answer by lwsshak3(11628)

Find the exact value:(without Calculator) Please show me how, i have an exam coming up! thank you in advance!
**
a) 2sin*2pie/3)-4tan*3pie/4
=2sin(2π/3)-4tan(3π/4)
=2*√3/2-4*-1=√3+4
..
b) 4tan*5pie/3+7cos*(-pie/4)
=4tan(5π/3)+7cos(-π/4)
=4*-√3+7*-√2/2
=-4√3-7√2/2